# Positive and Risky Message Assessment for Music Products

### v1

@@ 6 Ethical considerations and limitations
Social context concerns: The social context in which these labels are acquired is not always known, and there can be a lack of context in terms of why language is used and judged. For instance, many rap songs discuss the harms associated with isolation and substance abuse, yet such information might be misclassified due to rating provider bias or system bias. This may increase the likelihood of systematic bias or unintentionally promote racism. Future research should explore distinguishing between racist and reclaimed uses of slurs.
Data source concerns: Our rating data comes from CSM, a non-neutral organization. The regulations they use to recruit human experts to rate media products are unknown, and different experts rate different products, which may lead to inconsistencies. Transparency and accountability are not guaranteed, and subjectivity remains in the rating results. Future research and implementations should not rashly take ratings from data sources such as CSM as golden standard without careful assessment.
@@
Modality coverage: A technical limitation of this work is that it does not take other modalities of music products, such as melody, rhythm, and vocal performance, into account for predictions. We recognize the significance of these signals in conveying a song’s message. Additionally, we focused only on songs in English in this study, which could underrepresent lyrics written in other languages and from different cultural backgrounds. Future studies should include songs from various languages and cultures.
7 Conclusion
In this paper, we propose a novel task for the NLP community: predicting the strength of aspects in music where we cover objectionable aspects but also ones related to positive content. We leverage lyrics of music products to evaluate various dimensions of messages that they convey to the users. The proposed research problem and approach encourage and motivate further exploration of the music content assessment topic. The presented multi-task ordinality-enforcement model shows significant effectiveness for this task and inspires future research on practical problems with ordinal properties.
@@ Appendix A Implementation details
We list the implementation details of the proposed baseline methods. Sparse and dense document representation: Here we apply TF-IDF and Bag-of-Word-Vectors (GloVe Pennington et al. (2014) Average) as text vectorization approaches and linear models as baselines. TF-IDF method is also used in the previous work that classifies explicitness in lyrics Fell et al. (2020).
@@
Pretrained language model task fine-tuning: Contextualized representation from pretrained Transformer-based models has shown significant success on various NLP tasks. One popular variant, BERT Devlin et al. (2019), was also used in classifying explicitness in the previous workFell et al. (2020). We fine-tune a pretrained BERT on this task in the multi-class classification setting. The BERT model is adapted from HuggingFace with a maximum input length limitation of 512 tokens.
@@
A.1 Ranking-classification loss behavior

### v2

@@ 6.   Conclusion
Social context concerns: The social context in which these labels are acquired is not always known, and there can be a lack of context in terms of how language is used and judged. For instance, many rap songs discuss the harms associated with isolation and substance abuse, yet such information might be misclassified due to rating provider bias or system bias. This may increase the likelihood of systematic bias or unintentionally promote racism. Future research should explore distinguishing between racist and reclaimed uses of slurs as well as between mentions of risky subjects in a suggestive manner and those deemed more innocuous.
Data source concerns: Our rating data comes from CSM, a non-neutral organization. The regulations they use to recruit human experts to rate media products are unknown, and different experts rate different products, which may lead to inconsistencies. Transparency and accountability are not guaranteed, and subjectivity remains in the rating results. Future research and implementations should not rashly take ratings from data sources such as CSM as golden standards without careful assessment.
@@
Modality coverage: A technical limitation of this work is that it does not take other modalities of music products, such as melody, rhythm, and vocal performance, into account for predictions. We recognize the significance of these signals in conveying a song’s message. Additionally, we focused only on songs in English in this study, which means lyrics written in other languages and from different cultural backgrounds are absent from our study. Our work stands as an initial exploration of the feasibility of solving this task. We hope that future work can explore diversifying the dataset and exploring how the models behave in those cases.
Acknowledgements
We thank Common Sense Media for permitting us to use the expert ratings for research. We would like to thank the anonymous LREC-COLING reviewers for their feedback on this work.
7.   Bibliographical References
[…]
  • Aho and Ullman (1972) Alfred V. Aho and Jeffrey D. Ullman. 1972. The Theory of Parsing, Translation and Compiling, volume 1. Prentice-Hall, Englewood Cliffs, NJ.
  • American Academy of Pediatrics (1996) American Academy of Pediatrics. 1996. Impact of music lyrics and music videos on children and youth. Pediatrics, 98(6):1219–1221.
  • American Psychological Association (1983) American Psychological Association. 1983. Publications Manual. American Psychological Association, Washington, DC.
  • Ando and Zhang (2005) Rie Kubota Ando and Tong Zhang. 2005. A framework for learning predictive structures from multiple tasks and unlabeled data. Journal of Machine Learning Research, 6:1817–1853.
  • Andrew and Gao (2007) Galen Andrew and Jianfeng Gao. 2007. Scalable training of L1-regularized log-linear models. In Proceedings of the 24th International Conference on Machine Learning, pages 33–40.
  • BSI (1973a) BSI. 1973a. Natural Fibre Twines, 3rd edition. British Standards Institution, London. BS 2570.
  • BSI (1973b) BSI. 1973b. Natural fibre twines. BS 2570, British Standards Institution, London. 3rd. edn.
  • Castor and Pollux (1992) A. Castor and L. E. Pollux. 1992. The use of user modelling to guide inference and learning. Applied Intelligence, 2(1):37–53.
  • Chandra et al. (1981) Ashok K. Chandra, Dexter C. Kozen, and Larry J. Stockmeyer. 1981. Alternation. Journal of the Association for Computing Machinery, 28(1):114–133.
  • Chen et al. (2006) Meng-Jinn Chen, Brenda A Miller, Joel W Grube, and Elizabeth D Waiters. 2006. Music, substance use, and aggression. Journal of studies on alcohol, 67(3):373–381.
  • Chercheur (1994) J.L. Chercheur. 1994. Case-Based Reasoning, 2nd edition. Morgan Kaufman Publishers, San Mateo, CA.
  • Chin et al. (2018) Hyojin Chin, Jayong Kim, Yoonjong Kim, Jinseop Shin, and Mun Y Yi. 2018. Explicit content detection in music lyrics using machine learning. In 2018 IEEE International Conference on Big Data and Smart Computing (BigComp), pages 517–521. IEEE.
  • Chomsky (1973) N. Chomsky. 1973. Conditions on transformations. In A festschrift for Morris Halle, New York. Holt, Rinehart & Winston.
  • Devlin et al. (2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019. BERT: Pre-training of deep bidirectional transformers for language understanding. In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers), pages 4171–4186, Minneapolis, Minnesota. Association for Computational Linguistics.
  • Diaz and Marathe (2019) Raul Diaz and Amit Marathe. 2019. Soft labels for ordinal regression. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 4738–4747.
  • Eco (1990) Umberto Eco. 1990. The Limits of Interpretation. Indian University Press.
  • Fell et al. (2020) Michael Fell, Elena Cabrio, Elmahdi Korfed, Michel Buffa, and Fabien Gandon. 2020. Love me, love me, say (and write!) that you love me: Enriching the wasabi song corpus with lyrics annotations. In Proceedings of the 12th Language Resources and Evaluation Conference, pages 2138–2147.
  • Frank and Hall (2001) Eibe Frank and Mark Hall. 2001. A simple approach to ordinal classification. In European conference on machine learning, pages 145–156. Springer.
  • Gusfield (1997) Dan Gusfield. 1997. Algorithms on Strings, Trees and Sequences. Cambridge University Press, Cambridge, UK.
  • Hartmann (2022) Jochen Hartmann. 2022. Emotion english distilroberta-base. https://huggingface.co/j-hartmann/emotion-english-distilroberta-base/.
  • Hill (2011) Jennifer Ann Hill. 2011. Endangered childhoods: How consumerism is impacting child and youth identity. Media, Culture & Society, 33(3):347–362.
  • Hoel (1971a) Paul Gerhard Hoel. 1971a. Elementary Statistics, 3rd edition. Wiley series in probability and mathematical statistics. Wiley, New York, Chichester. ISBN 0 471 40300.
  • Hoel (1971b) Paul Gerhard Hoel. 1971b. Elementary Statistics, 3rd edition, Wiley series in probability and mathematical statistics, pages 19–33. Wiley, New York, Chichester. ISBN 0 471 40300.
  • Jespersen (1922) Otto Jespersen. 1922. Language: Its Nature, Development, and Origin. Allen and Unwin.
  • Kim (2014) Yoon Kim. 2014. Convolutional neural networks for sentence classification. In Proceedings of the 2014 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 1746–1751, Doha, Qatar. Association for Computational Linguistics.
  • Kolbe et al. (1993) Lloyd J Kolbe, Laura Kann, and Janet L Collins. 1993. Overview of the youth risk behavior surveillance system. Public Health Reports, 108(Suppl 1):2.
  • Lai et al. (2015) Siwei Lai, Liheng Xu, Kang Liu, and Jun Zhao. 2015. Recurrent convolutional neural networks for text classification. Proceedings of the AAAI Conference on Artificial Intelligence, 29(1).
  • Martinez et al. (2020) Victor Martinez, Krishna Somandepalli, Yalda Tehranian-Uhls, and Shrikanth Narayanan. 2020. Joint estimation and analysis of risk behavior ratings in movie scripts. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 4780–4790, Online. Association for Computational Linguistics.
  • Martinez et al. (2019) Victor R Martinez, Krishna Somandepalli, Karan Singla, Anil Ramakrishna, Yalda T Uhls, and Shrikanth Narayanan. 2019. Violence rating prediction from movie scripts. In Proceedings of the AAAI Conference on Artificial Intelligence, volume 33, pages 671–678.
  • OpenAI (2022) OpenAI. 2022. Introducing chatgpt. Accessed: 2023-10-14.
  • Ouyang et al. (2022) Long Ouyang, Jeffrey Wu, Xu Jiang, Diogo Almeida, Carroll Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, et al. 2022. Training language models to follow instructions with human feedback. Advances in Neural Information Processing Systems, 35:27730–27744.
  • Pennington et al. (2014) Jeffrey Pennington, Richard Socher, and Christopher Manning. 2014. GloVe: Global vectors for word representation. In Proceedings of the 2014 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 1532–1543, Doha, Qatar. Association for Computational Linguistics.
  • Primack et al. (2008) Brian A Primack, Madeline A Dalton, Mary V Carroll, Aaron A Agarwal, and Michael J Fine. 2008. Content analysis of tobacco, alcohol, and other drugs in popular music. Archives of pediatrics & adolescent medicine, 162(2):169–175.
  • Primack et al. (2009) Brian A Primack, Erika L Douglas, Michael J Fine, and Madeline A Dalton. 2009. Exposure to sexual lyrics and sexual experience among urban adolescents. American journal of preventive medicine, 36(4):317–323.
  • Rasooli and Tetreault (2015) Mohammad Sadegh Rasooli and Joel R. Tetreault. 2015. Yara parser: A fast and accurate dependency parser. Computing Research Repository, arXiv:1503.06733. Version 2.
  • Reimers and Gurevych (2019) Nils Reimers and Iryna Gurevych. 2019. Sentence-BERT: Sentence embeddings using Siamese BERT-networks. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), pages 3982–3992, Hong Kong, China. Association for Computational Linguistics.
  • Rospocher and Eksir (2023) Marco Rospocher and Samaneh Eksir. 2023. Assessing fine-grained explicitness of song lyrics. Information, 14(3).
  • Ruder (2017) Sebastian Ruder. 2017. An overview of multi-task learning in deep neural networks. arXiv preprint arXiv:1706.05098.
  • Samghabadi et al. (2020) Niloofar Safi Samghabadi, Afsheen Hatami, Mahsa Shafaei, Sudipta Kar, and Thamar Solorio. 2020. Attending the emotions to detect online abusive language. In Proceedings of the Fourth Workshop on Online Abuse and Harms, pages 79–88.
  • Shafaei et al. (2020) Mahsa Shafaei, Niloofar Safi Samghabadi, Sudipta Kar, and Thamar Solorio. 2020. Age suitability rating: Predicting the MPAA rating based on movie dialogues. In Proceedings of the 12th Language Resources and Evaluation Conference, pages 1327–1335, Marseille, France. European Language Resources Association.
  • Singer et al. (1954–58) Charles Joseph Singer, E. J. Holmyard, and A. R. Hall, editors. 1954–58. A history of technology. Oxford University Press, London. 5 vol.
  • Strötgen and Gertz (2012) Jannik Strötgen and Michael Gertz. 2012. Temporal tagging on different domains: Challenges, strategies, and gold standards. In Proceedings of the Eight International Conference on Language Resources and Evaluation (LREC’12), pages 3746–3753, Istanbul, Turkey. European Language Resource Association (ELRA).
  • Superman et al. (2000) S. Superman, B. Batman, C. Catwoman, and S. Spiderman. 2000. Superheroes experiences with books, 20th edition. The Phantom Editors Associates, Gotham City.
  • Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-thought prompting elicits reasoning in large language models. Advances in Neural Information Processing Systems, 35:24824–24837.
  • Zhang et al. (2021) Yigeng Zhang, Mahsa Shafaei, Fabio Gonzalez, and Thamar Solorio. 2021. From none to severe: Predicting severity in movie scripts. In Findings of the Association for Computational Linguistics: EMNLP 2021, pages 3951–3956, Punta Cana, Dominican Republic. Association for Computational Linguistics.
@@ Appendix A Implementation details
We list the implementation details of the proposed baseline methods.
Sparse and dense document representation: Here we apply TF-IDF and Bag-of-Word-Vectors (GloVe Pennington et al. (2014) Average) as text vectorization approaches and linear models as baselines. TF-IDF method is also used in the previous work that classifies explicitness in lyrics Fell et al. (2020).
@@
Pretrained language model task fine-tuning: Contextualized representation from pretrained Transformer-based models have shown significant success on various NLP tasks. One popular variant, BERT Devlin et al. (2019), was also used in classifying explicitness in the previous workFell et al. (2020). We fine-tune a pretrained BERT on this task in the multi-class classification setting. The BERT model is adapted from HuggingFace with a maximum input length limitation of 512 tokens.
@@
A.1.   Ranking-classification loss behavior
@@
A.2.   Prompt used for LLM evaluation
Simple context: Please assess the lyrics of the song given the following aspects: The lyrics: <Full Lyrics> Positive Messages; Violence & Scariness; Sex, Romance & Nudity; Drinking, Drugs & Smoking; Products & Purchases (refer to Consumerism).
Please rate the presence of each aspect on a scale of 0 to 2, where 0 indicates ’low’, 1 indicates ’medium’, and 2 indicates ’high’. Provide your ratings strictly in the JSON format as shown in the example below and make sure no extra content: Example: {"Positive Messages": 1, "Violence & Scariness": 0, "Sex, Romance & Nudity": 1, "Drinking, Drugs & Smoking": 2, "Products & Purchases": 1}.
Rich context: Please assess the lyrics of the song given the following aspects: The lyrics: <Full Lyrics> Positive Messages: <Full aspect description from CSM>; Violence & Scariness: <Full aspect description from CSM>; Sex, Romance & Nudity: <Full aspect description from CSM>; Drinking, Drugs & Smoking: <Full aspect description from CSM>; Products & Purchases (refer to Consumerism): <Full aspect description from CSM>.
Please rate the presence of each aspect on a scale of 0 to 2, where 0 indicates ’low’, 1 indicates ’medium’, and 2 indicates ’high’. Provide your ratings strictly in the JSON format as shown in the example below and make sure no extra content: Example: {"Positive Messages": 1, "Violence & Scariness": 0, "Sex, Romance & Nudity": 1, "Drinking, Drugs & Smoking": 2, "Products & Purchases": 1}.
Chain-of-Thought (CoT) prompt:
(Keep context part the same).
Please rate the presence of each aspect on a scale of 0 to 2, where 0 indicates ’low’, 1 indicates ’medium’, and 2 indicates ’high’. Let’s think step-by-step to analyze the lyrics and then provide your ratings in the JSON format. Here is an example output: This song promotes ... and the song has strong ... It depicts ..., so it implies ... (your chain-of-thought) ... Therefore, we reach the final assessment result: {"Positive Messages": 1, "Violence & Scariness": 0, "Sex, Romance & Nudity": 1, "Drinking, Drugs & Smoking": 2, "Products & Purchases": 1}. Now it is your turn: